6. An airplane whose air speed is 650km/h is supposed to fly in straight path 35° north of
east. But a steady 95 km/h wind is blowing from north. In what direction should the
plane head and what is the plane resultant speed?
1
Expert's answer
2020-05-12T09:59:10-0400

This problems is related to the notion of addition of vectors.

The resultant velocity "\vec{v}_{rel} = \vec{v}_{wind} + \vec{v}_{plane}." We know that the wind counteracts to the movement of

the plane, so the resultant velocity will be smaller than the initial velocity "v_{plane}" . Also, we conclude that

the angle in question should be larger than "35^\circ" .


The angle between "\vec{v}_{wind}" and "\vec{v}_{res}" should be "\alpha = 90^\circ + 35^\circ = 125^\circ." We may calculate the angle "\beta" between "\vec{v}_{res}" and "\vec{v}_{plane}" using the law of sines

"\dfrac{v_{plane}}{\sin\alpha} = \dfrac{v_{wind}}{\sin\beta}."

Therefore, "\sin\beta = \dfrac{v_{wind}}{v_{plane}}\sin\alpha = \dfrac{95}{650} \sin125^\circ \approx 0.12, \quad \beta = 6.9^\circ." The third angle "\gamma" between "v_{wind}" and "v_{plane}" is "\gamma = 180^\circ - \alpha-\beta = 180^\circ - 125^\circ-6.9^\circ= 48.1^\circ."


Therefore, we may calculate "v_{res}" using the law of cosines

"v_{res}^2 = v_{wind}^2 + v_{plane}^2 - 2v_{wind}v_{plane}\cos\gamma."

"v_{res}\approx 590\,km/h."


The angle in question is "90^\circ - \gamma = 90^\circ-48.1 = 41.9^\circ" north of east.